OpenStudy (anonymous):

Will fan and award :) The radius of a cone-shaped tank is 4 feet less than its height. If the height of the tank is (x - 3) feet, the expression below shows the volume of the tank. 1 over 3π (x - 7)2 (x - 3) What does the factor π(x - 7)2 (x - 3) represent?

OpenStudy (anonymous):

The area of the base of the tank The area of the curved sides of the tank The volume of 3 of the same cone-shaped tanks The volume of 6 of the same cone-shaped tanks

OpenStudy (bibby):

it might help if you rewrite the formula. 1/3π (x - 7)2 (x - 3) where π (x - 7)2 (x - 3) = abc (just a placeholder name) so you have volume = 1/3 * abc what is abc?

OpenStudy (anonymous):

Im stilll confused, is there any other way you could put it? would abc be the answer im looking for?

OpenStudy (bibby):

\[\frac{ 1 }{3 } * abc\]= volume of the tank. what is abc? (hint it's 3 times the above quantity)

OpenStudy (anonymous):

its 3 times 1/3 ?

OpenStudy (bibby):

How do you get \[\frac{1}{3}x -> x\]

OpenStudy (anonymous):

aha i get it , thanks

OpenStudy (bibby):

lol should have used x instead of abc

OpenStudy (anonymous):

yes , but either way , was still fine :)